Find an explicit solution of the given initial- value problem

square root of 1-y^2 dx - square root 1-x^2 dy = 0, y(0)= 1/2

y = _________________
1
Expert's answer
2020-07-21T18:44:59-0400

As,

"\sqrt{1-y^2}dx-\sqrt{1-x^2}dy=0"

Divide both side of the above equation by "\sqrt{(1-y^2)(1-x^2)}" ,thus


"\dfrac{dx}{\sqrt{1-x^2}}-\dfrac{dy}{\sqrt{1-y^2}}=0\\
\implies \int\dfrac{dx}{\sqrt{1-x^2}}-\int\dfrac{dy}{\sqrt{1-y^2}}=c\\
\implies\sin^{-1}(x)-\sin^{-1}(y)=c"

Now, "y(0)=1\/2" ,thus

"\sin^{-1}(0)-\sin^{-1}(1/2)=c"

Thus, particular solution is

"\sin^{-1}(x)-\sin^{-1}(y)=-\frac{\pi}{6}"

Thus,

"y=\sin(\sin^{-1}(x)+\frac{\pi}{6})"